Bryce insists that(2+x)/8 is equivalent to (1+x)/4. What mistake do you think is being made and how could you demonstrate to Bryce that the two expressions are not equal?
Answer by fcabanski(1391)

He's treating 2x+8 as 2x*8, dividing both the 2 and 4 by 2.


Tell him to pick a number for x. Let's pick 1.


(2+1)/8 = 3/8 and (1+1)/4 = 2/4. They are not equal.